OpenStudy (anonymous):

Kelly cannot remember the formula for the sum of a finite arithmetic sequence, so she is going to try to derive it. Her problem is to find the sum of the numbers 1 to 500. What should she do first?

OpenStudy (anonymous):

1+2+3+....+500=X Use the summation of series

OpenStudy (anonymous):

what a weird question, but i am going to guess at the answer

OpenStudy (anonymous):

add forwards \[1+2+3+..+498+499+500\] and backwards \[500+499+498+...+3+2+1\] get twice the sum as \[501+501+501+...+501+501+501=500\times 501\] and since this is twice the sum you want, divide by 2 and get \[\frac{500\times 501}{2}\]